OpenStudy (anonymous):

Help: What is the equation of a line that passes through (4, 7) and has a slope of −14? y=x4−74 y=−x4+74 y=−x4+8 y=x4+8

OpenStudy (mrnood):

I think you have typed this wrong is the slope -4 (or -14) answers would normally be something like y=-14x +10 why are they written y = x4 ....?

OpenStudy (anonymous):

What is the equation of a line that passes through (4, 7) and has a slope of −1/4? y=x/4−7/4 y=−x/4+7/4 y=−x/4+8 y=x/4+8

OpenStudy (anonymous):

sorry they where all fractions

OpenStudy (mrnood):

ok - that is better so - a straight line has an equation y=mx+b where m and b are constant values m is the slope SO to start with: you know the slope is -1/4 which 2 answers have a slope of -1/4 ?

OpenStudy (anonymous):

a and b?

OpenStudy (mrnood):

 which have a value of -1/4 in front of the x?